OpenStudy (anonymous):

2 minor league baseball players got a total of 314 hits. Washington had 4 more hits than Sanchez. Find the number of hits for each player. Sorry but I hate these types.

OpenStudy (anonymous):

Let's say that Washington got x number of hits and Sanchez got y number of hits. So, x+y=314, because the total is 314 hits. We know that Washington (x hits) got 4 more hits than Sanchez (y hits), so really, x=y+4. If we substitute x=y+4 into x+y=314, we have (y+4)+y=314, or 2y=310. Divide by two and you have y =155. So Sanchez got 155 hits, and Washington got 159.

OpenStudy (anonymous):

w=s+4 w+s=314 s+4+s=314 2s = 310 s= 155 w=s+4 w=159

OpenStudy (anonymous):

you all are really lifesavers. it has been 40 yrs since i had basic math. this is all new to me. thanks again

OpenStudy (anonymous):

Your welcome!